2FeS2(g)+1/2O2(g)−→Fe2O3+4SO2(g). TakingtheenthalpyofformationofFeS2(g),Fe2O3 andSO2(g) to be -180, -824 and -297 kJmol−1. If the starting material FeS2 is 0.1, what is the value of enthalpy of reaction and also if the C −v of the calorimeter is 10 kJK−1, what will be the change in the temperature.


1
Expert's answer
2021-02-28T06:20:53-0500

2FeS2 (g) + 1/2 O2 "\to" Fe2O3 + 4 SO2 (g)

-180 _____0__________ - 824____ - 297

"\Delta H" = - 824 - 4(-297) - 2(-180)

= -824 -1188 +360 KJ/mol

= - 1652 KJ/mol

"\Delta H" = - n c "\Delta T"

- 1652 = - 0.1 × 10 × "\Delta T"

"\Delta T" = 165.2